ABSOLUTE INTENSITIES FOR THE O_{2} 1.27 μm CONTINUUM ABSORPTION

\underline{B. MATÉ}, C. LUGEZ, G.T. FRASER and W.J. LAFFERTY, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

Collision-induced absorption coefficients for the 1.27 μm band of O_{2} have been measured at a resolution of 0.5 cm^{−1} and an optical pathlength of L = 84 m using a Fourier-transform spectrometer and 2-m long White-type multipass absorption cell. Spectra were recorded for sample densities, ρ, from 1 to 10 times that of a ideal gas under standard conditions (T = 273.15 K and P = 101.325 kPa), i.e., 1 to 10 amagats, at temperatures of 253 K, 273 K, and 296 K, for pure O_{2} and O_{2}/N_{2} mixtures. After removing the contributions from the sharp lines of the v = 0 ← 0 component of the O_{2} a^{1}Δ_{g} ← X^{3}Σ_{g}^{−} band, which overlaps the continuum band, the integrated band strength per unit pathlength, S ≡ S_{O_{2}−O_{2}} ρ_{O_{2}}^{2} + S_{O_{2}−N_{2}} ρ_{N_{2}} ρ_{O_{2}}, has been determined for several values of the densities, ρ_{O_{2}} and ρ_{N_{2}}, to give values for S_{O_{2}−O_{2}} and S_{O_{2}−N_{2}}. At 296 K we find S_{O_{2}−O_{2}} = 4.847(22) × 10^{−43} cm^{−2} (molecule/cm^{3})^{−2} [3.499(16) × 10^{−4} cm^{−2} amagat^{−2}] and S_{O_{2}−N_{2}} = 0.941(50) × 10^{−43} cm^{−2} (molecule/cm^{3})^{−2} [0.679(36) × 10^{−4} cm^{−2} amagat^{−2}]. Here and elsewhere, Type A expanded uncertainties are given with a coverage factor k = 2. The S_{O_{2}−O_{2}} coefficient is in reasonable agreement with the previous measurements of Cho et al. [C.W. Cho, E.J. Allin, and H.L. Welsh, Can.J.P hys. 41, 1991-2002 (1963)], however our value of S_{O_{2}−N_{2}} is a factor of 2.6 times greater than their results. The derived air coefficient, S_{O_{2}−air}, is 37 % greater than the value determined by Mlawer et al. [E.J. Mlawer, S.A. Clough, P.D. Brown, T.M. Stephen, J.C. Landry, A. Goldman, and F.J. Murcray, J.Geophys.Res. 103, 3859-3863 (1998)] from an atmospheric measurement, which has been corrected for the most recent value for the absorption coefficients for the overlapping O_{2} a^{1}Δ_{g} ← X^{3}Σ_{g}^{−} band. The binary collision coefficients are available as a function of frequency for use in atmospheric modeling.